OpenStudy (pphalke):

Evaluate 4! which formula would I use?

jimthompson5910 (jim_thompson5910):

The exclamation point after the number means `factorial` You start with the given number, and multiply ALL of the whole numbers smaller than it til you get to 1 Examples: 7! = 7*6*5*4*3*2*1 = 5040 9! = 9*8*7*6*5*4*3*2*1 = 362,880 5! = 5*4*3*2*1 = 120